OpenStudy (anonymous):

Find the exact value of cos^-1(- sqrt3/2)

OpenStudy (anonymous):

look on the unit circle and find the place in the upper half where the first coordinate is \(-\frac{\sqrt{3}}{2}\) the angle you see is your answer

OpenStudy (anonymous):

I came up with (5π)/(6), can you help show me the steps to get to that answer so that I will better understand. Thank You

OpenStudy (anonymous):

\(\theta = \cos^{-1}h\) means "find the angle \(\theta\) who's \(x\) value on the unit circle equals \(h\)".

OpenStudy (anonymous):

Also, like @satellite73 said, \(\cos^{-1}\) has a domain of \([-1,1]\) and a range of \([0,\pi]\), meaning you can only use angles on the top half of the circle.